OpenStudy (anonymous):

please help 2x+5y=-24 3x-5y=14 at what 2 points do thease lines intersect

OpenStudy (bahrom7893):

set them equal to each other..

OpenStudy (bahrom7893):

sorry meant solve both for y and x

OpenStudy (bahrom7893):

2x + 5y = -24 2x = -24 - 5y x = -12 - (5/2)y

OpenStudy (bahrom7893):

plug that value into the previous one..

OpenStudy (bahrom7893):

3(-12 - (5/2)y) - 5y = 14

OpenStudy (anonymous):

then do disributive propertys right

OpenStudy (anonymous):

The easiest approach to this kind of problem is to solve each equation for y. You should be able to do that. Then set the two non-y halves equal to each other. Example: y = 2x + 1 and y = 3x - 5 Set 2x + 1 = 3x - 5 Solve for x. Finally, plug the x into either of the y equations to get y. Done.

OpenStudy (bahrom7893):

lol Lbickford's approach is easier!

OpenStudy (anonymous):

haha ok i understand now thanks both of u guys

OpenStudy (bahrom7893):

can you click on "Become a fan" next to both of our names? Thanks =)